Write an equation in slope intercept form, y = mx + b, of the line that has a slope of –2/3, and contains the point, (6, –5).
Answer by checkley71(8403)

Y=-2/3X+b
NOW SUBSTITUTE THE X & Y VALUES OF THE POINT THUS & SOLVE FOR THE Y INTERCEPT(b)
-5=-2/3*6+b
-5=-4+b
b=-5+4
b=-1
THUS THE EQUATION IS
Y=-2/3X-1